### v2

# Passive defect driven morphogenesis in nematic membranes

Abstract
Orientational order is a common feature of many biological and synthetic materials. Topological defects are discontinuities in this order that are often coupled to geometric features of the materials. Here, we study the equilibrium shapes of fluid membranes featuring a $+1$ topological defect as a model for morphogenesis. We show, through simulation and analytic calculation, that the membrane can spontaneously deform toward a conical shape with a defect at its apex. We show that the relative stability of the deformation is controlled by the balance of the elastic parameters. When boundary constraints are introduced, we observe three distinct modes of deformation. These deformation modes take advantage of the way in which splay, twist and bend distortions of the director field can be exchanged on a curved surface. Finally we demonstrate inverted solutions. Our findings demonstrate a mechanism for passive defect driven morphogenesis as well as the fusion of $+1/2$ topological defect pairs on deformable surfaces.
The study of shape shifting materials is often inspired by biological systems and has a broad range of applications. For example, deformable solids that have a prescribed orientational field, such as elastomers [1, 2, 3] or inflatable structures [4, 5], can achieve families of morphologies by differential growth [6, 7, 8, 9, 10] and have applications in soft robotics[11]. Here, we focus on fluid membranes with nematic order. Unlike the previous cases, these membranes can, in addition, flow, remodel and adapt both their orientational field and shape.
Fluid membranes achieve their shapes by minimizing a free-energy subjected to physical constraints [12, 13, 14, 15]. When orientational order is included, the stresses generated by topological defects can render a flat surface unstable to out-of-plane perturbations [16]. In addition, the geometrical properties of the surface, such as the extrinsic curvature, can in turn influence the dynamics of the orientational field and its equilibrium configurations [17, 18, 19, 20, 21, 22, 23, 24]. Indeed, for prescribed geometries, couplings between the nematic field and the intrinsic or the extrinsic geometry have been shown to induce symmetry-breaking of orientational configurations, or influence topological defect dynamics, [17, 18, 25, 26, 27, 19]. Recent studies extended these works by, for instance, including out-of-equilibrium processes, such as active stresses or active torques, or varying surface geometry and topology [28, 29, 30, 31, 32, 33, 34, 35, 36].
We study rotationally symmetric systems, which therefore feature a $+1$ topological defect at their center; $+1$ topological defects have been associated with geometrical changes in natural and synthetic systems [37, 38, 39, 40]. We use cylindrical coordinates to define these shapes, where $r$ is the radial coordinate, $\theta$ is the azimuthal coordinate and $\zeta$ is the axial coordinate, Fig. 1a. The outer circular boundary is placed at $r=R$; without loss of generality, we set the vertical offset by $\zeta(R)=0$ and $R=1$. The orientation of the nematic field is described by a director field $\hat{n}$ that represents the averaged local orientation on the surface. We consider that the director field is tangential to the surface. In addition, we consider that the system is deep into the nematic phase and impose $|\hat{n}|=1$. This allows the director field to be defined by a scalar phase $\psi(r)$, which corresponds to the angle between the director field and the curvilinear radial direction. The cases $\psi=0$ corresponds to the aster, $20<\psi<\pi/$ 2 to a spiral, and $\psi=\pi/2$ to the vortex.
The two-dimensional free-energy of a fluid membrane with nematic order is given by
$\mathcal{F}$ $=$ $\int_{\mathcal{A}}\Big{\{}k_{B}{\cal H}^{2}+\sigma+k_{1}(\nabla\cdot\hat{n})^{2}$ (1)
$+$ $k_{2}(\hat{n}\cdot(\nabla\times\hat{n}))^{2}+k_{3}(\hat{n}\times(\nabla\times\hat{n}))^{2}\Big{\}}\textrm{d}a.$
The first term is the bending energy with mean curvature ${\cal H}$, and the second term represents surface tension. We disregard anisotropies in bending energy dependent on $\hat{n}$. The other terms are the Frank free-energy associated respectively with splay, twist and bend distortions of the director field $\hat{n}$ [41, 42]. The corresponding elastic coefficients are: the bending rigidity $k_{B}$, the surface tension $\sigma$, and the reduced Frank constants $k_{1}$, $k_{2}$, and $k_{3}$ that are proportional to the membrane thickness. Because the director field is parallel to the surface, the effects of the saddle-splay distortions with elastic constant $k_{24}$ can be absorbed in a redefinition of the other Frank constants, [18]. As shown in Refs. [17, 18] and subsequently extended for more general cases in [43], the thin-film limit of the Frank free-energy results in contributions that couple the director field with both the intrinsic and the extrinsic geometry. In our case, Eq. (S2) takes the form derived in Refs. [17, 18] and the expression of the free-energy (S2) for the special case of a surface of revolution is derived in Sec. 1.1 [44]. Other descriptions for fluid surfaces with orientational order neglected the couplings with the extrinsic geometry [16, 26, 25]. For a discussion on the thin-film approximation of liquid crystals, we refer to Ref. [43].
We combine analytical and numerical approaches to study equilibrium configurations of the free-energy (S2). The former approach is restricted to director fields with a uniform phase $\psi$, see Sec. 1 and 2 in [44]. The latter approach discretizes the functions $\zeta(r)$ and $\psi(r)$, Fig. 1a, and uses a Monte-Carlo algorithm, see Sec. 3 [44].
Figure 1: (a) Geometrical setup. The red line shows the height profile, $\zeta(r)$, and black dashes show the orientation of the director, $\psi(r)$. (b) The height of a membrane featuring a single aster defect with equal Frank constants $k_{1}\!=\!k_{2}\!=\!k_{3}\!=\!k/3$ as a function of $k$, surface tension $\sigma$ and bending rigidity $k_{B}$ with the constraint $\sigma+k_{B}+k=1$. The magenta line corresponds to Eq. S5. From this point forward we fix $\sigma=k_{B}=1/10$ and $k_{1}+k_{2}+k_{3}=1$ in all results unless otherwise stated. (c) Height of a membrane as a function of the prescribed uniform phase, $\psi$, and ratio between the splay and bend coefficients, ($k_{2}=0$). The green line corresponds to the predicted transition to conical surfaces, see SI. (d)&(e) Examples of equilibrium surfaces with parameters given by: white star ($k_{1}=k_{3}$, $\psi=0$), red star ($k_{3}=1$, $\psi=\pi/8$). (d) An aster on a conical surface, and (e) a spiral on a concave surface. Blue (red) regions: negative (positive) Gaussian curvature.
We first consider the case where all the reduced Frank constants are equal, i.e. the one-constant approximation $k_{1}\!=\!k_{2}\!=\!k_{3}\!=\!k/3$. The energy scale is set by the constraint $k\!+\!k_{B}\!+\!\sigma\!=\!1$. Fig. 1b shows the height at the center $\zeta(0)$ of the equilibrium states that were found for varying elastic parameters. When $k_{B}$ or $\sigma$ dominate, the minimal states are flat discs which minimize surface area and curvature. Because the Frank constants are equal, $k_{1}\!=\!k_{3}$, all director configurations with a constant phase $\psi$ have equal energy, and thus are minimal states, [42]. However, when $k$ dominates, the minimal state changes to an approximately conical surface with phase $\psi=0$, Fig. 1d. The morphological transition occurs via a symmetry-breaking process during which the director adopts an aster configuration.
To better understand this spontaneous transition from flat to conical surfaces, we derived analytically the normal force balance equation for a surface of revolution with an embedded director field with uniform phase $\psi$, see Sec. 1.2 in [44] (see [27] for a general case). In the special case $\sigma=0$, this nonlinear ODE has a set of exact non-trivial solutions corresponding to conical surfaces of varying heights; this is valid even when the one-constant approximation is relaxed, see Sec. 1.3 in [44]. This gives us a subset of shapes described by $\zeta(r)=\pm m(r-1)$, with $m$ the vertical distance of the tip of the cone from the base, that allow us to move continuously from a flat disk to a cone of varying height. Minimizing Eq. (S2) with respect to the height $m$ and for a fix phase $\psi$, one obtains the non-trivial condition
$m^{2}=\frac{A(1-2\sin(\psi)^{2})-2-B}{1+A\sin{(\psi)}^{2}+B},$ (2)
where $A=4k/3k_{B}$ and $B=(4\sigma/k_{B})(1-\Delta^{2})/\log(1/\Delta)$ are dimensionless parameters and $\Delta$ is a dimensionless cut-off lengthscale. The logarithmic divergence at $\Delta=0$ is commonly found in the context of topological defects [42] and also describes the divergence in curvature at the tip of the conical surface. We explore the effect of $\Delta$ on our results in Sec. 2 in [44]. Eq. (S3) leads to the existence condition for conical shapes $A(1-2\sin(\psi)^{2})>2+B$. As observed, Eq. (S3) shows that as $k/k_{B}$ or $k/\sigma$ increase, the minimal surface varies from a flat disc ($m=0$) to a cylinder ($m\rightarrow\infty$). The total free-energy associated with this minimal surface reads
$\frac{{\cal F}_{c}}{\pi k_{B}}=\sqrt{(1+A\sin{(\psi)}^{2}+B)(A\cos{(\psi)}^{2}-1)}\log(1/\Delta).$ (3)
In addition, Eq. (S4) shows that the minimal phase corresponds to $\psi=0$, even when the Frank constants $k_{1}\!=\!k_{3}$ are equal. The selection mechanism for the phase arises from the coupling between the director field and the extrinsic curvature, which tend to align the director field with the minimal principal curvature [18]. In the case of conical surfaces, the aster is favoured because it features only splay distortions, and vanishing twist and bend distortions. Furthermore, the total Frank free-energy for the aster decreases as the height of the cone increases, leading to the spontaneous out-of-plane deformation of a surface. This is balanced by the increased curvature and area of the surface. The threshold for a flat disc with an aster topological defect to become unstable is set when its energy (i.e. $2\pi k/3\log(1/\Delta)+\sigma\pi(1-\Delta^{2})$) equals the energy (S4) for $\psi=0$, that is when
$k/3=\frac{k_{B}}{2}+\sigma\frac{(1-\Delta^{2})}{\log(1/\Delta),}$ (4)
shown with the magenta line, Fig. 1b. This condition equals the existence condition of conical shapes for $\sigma=0$, Sec. 1.3 in [44]. A similar approach is used with varying elastic coefficients and defect phase to generate the green line in Fig. 1c.
According to Eq. S3, the height of the surface is also controlled by the phase of the defect. For equal Frank constants, the existence condition for conical shapes can be satisfied when $\psi<\pi/4$. Beyond this point, the increased energy associated with twist and bend distortions of the director field on a conical surface is too great and a flat surface is favoured. To further explore this effect, we numerically studied the equilibrium shapes of surfaces with a prescribed uniform phase, $\psi$, and a varying ratio of bend and splay elastic constants, $k_{1}/(k_{1}+k_{3})$, Fig. 1c. In general, we find that vortices cannot deform a surface. In most cases, the aster generates the maximal out-of-plane deformation, except when $k_{3}\gg k_{1}$ and the maximal deformation occurs near $\psi\sim\pi/8$, see Fig. 1e. This result is also predicted by the conical surface approximation, which is given by the green line in Fig. 1c.
Figure 2: (a) Height at the center and (b) phase at the boundary of the surface as a function of the three Frank constants with the constraint $k_{1}+k_{2}+k_{3}=1$. Magenta line is Eq. (S5) and red line is Eq. (S6). (c)&(d) Example surfaces with parameters: green star ($k_{1}=1/2$, $k_{2}=1/10$, $k_{3}=4/10$) blue star ($k_{1}=13/20$, $k_{2}=1/10$, $k_{3}=1/4$). (c) shows a conical aster surface while (d) features a concave surface with an aster at the center and spiral at the boundary. Blue (red) regions: negative (positive) Gaussian curvature. In all panels, $k_{B}=\sigma=1/10$.
To further explore these results, we now focus on the effects of varying all three Frank elastic constants; we also relax the constraint that $\psi$ is constant. Both $\zeta^{\prime}$ and $\psi$ are only constrained by the rotational symmetry at the boundary. Figs. 2a and 2b show respectively the height at the center, $\zeta(0)$, and the phase at the outer boundary, $\psi(R)$, of the energy minimizing states for varying Frank coefficients. We found that the height of the deformed membrane is determined by the magnitude of $k_{1}$. Two transitions from flat to deformed surfaces were identified, which are primarily dependent on the relative values of $k_{1}$ and $k_{3}$. The transition from a flat to a deformed surface with an aster is determined by the threshold (S5) with $k_{1}=k/3$ (magenta line in Figs. 2a-b). A flat disc with a vortex is linearly stable to out-of-plane deformations, see Sec. 1.4 in [44]. The red line in Fig. 2a-b can be found by comparing the energy of a flat disc with a vortex (i.e. $2\pi k_{3}\log(1/\Delta)+\pi\sigma(1-\Delta^{2})$) to the energy (S4) for $\psi=0$, that is when
$\frac{k_{1}}{k_{B}}=\frac{(k_{3}/k_{B}+B/4)^{2}}{(1+B)}+\frac{1}{4}.$ (5)
On a flat surface, if $k_{1}\!>\!k_{3}$ ($k_{1}\!<\!k_{3}$), the director field assumes a vortex (aster) configuration [42]. On a curved surface, however, an aster can be energetically favoured for values of $k_{1}>k_{3}$; indeed all deformed surfaces here feature an aster at their core. In most cases, this is a conical aster deformation with constant phase, Fig. 2c. However, when $k_{2}\ll k_{3}<k_{1}$ we observe a new state with a spatially varying phase which features a conical aster close to the core combined with a negative Gaussian curvature spiral region near the boundary, Fig. 2d. The spiral region reduces splay at the cost of additional bend, however the bend is then exchanged for twist on the negative Gaussian curvature surface when $\psi$ has an intermediate value, which significantly reduces the free-energy density when $k_{2}$ is small.
Figure 3: (a) Height and (b) phase of the director field at the center of the membrane as a function of the phase at the boundary, $\psi(R)=\psi_{R}$, and the ratio between bend and splay coefficients ($k_{2}=1/3$) for surfaces with fix boundaries $\zeta^{\prime}(R)=0$. (c-e) Example of the three modes of out-of-plane deformation with fix boundary conditions: (c) pointy with an aster, (d) domed with an aster-to-vortex transition and (e) pointy with a vortex-to-aster transition. The states in (c-e) correspond to the black ($k_{1}=1/2$, $k_{3}=1/6$, $\psi_{R}=0$), turquoise ($k_{1}=2/3$, $k_{3}=0$, $\psi_{R}=0$) and pink ($k_{1}=0$, $k_{3}=2/3$, $\psi_{R}=\pi/2$) stars in (a) respectively. Blue (red) regions: negative (positive) Gaussian curvature. In all panels, $k_{B}=\sigma=1/10$.
Now, we consider the system under boundary conditions that might be found in biological systems. Hence, we fix $\psi(R)\!=\!\psi_{R}$ as the phase at the boundary of the membrane and fix $\zeta^{\prime}(R)\!=\!0$, implying that the surface must be flat at its boundary  [^1] Note that the boundary condition $\zeta^{\prime}(R)\!=\!0$ enforces that the total Gaussian curvature, including the tip, is zero.. Figs. 3a&b show the height and the phase at the center of the surface for a full range of the bend and splay constants ($k_{2}=1/3$) and the boundary phase $\psi_{R}$. On a flat surface, the value of $k_{1}/(k_{1}+k_{3})$ controls the phase at the core of the topological defect. By varying $\psi_{R}$, we can explore regimes where the preferred phase at the center is frustrated with the boundary. We observe five distinct states. First, when the boundary phase $\psi_{R}$ is compatible with the dominant elastic coefficient, we observe flat asters and flat vortices. In addition to this there are three deformed configurations.
When $k_{1}\!\approx\!k_{3}$ and $\psi_{R}\!\approx\!0$ we find pointy surfaces featuring an aster, where the core and boundary of the defect are broadly in phase, see Fig. 3c. As in Fig. 2, the deformation here reduces splay distortions and hence stabilizes an aster, even in some cases when $k_{1}>k_{3}$.
Figure 4: (a) Height at the apex as a function of the magnitude of $k_{2}$ for the three modes of out-of-plane deformation in Fig. 3c-e. The values of $k_{1}$, $k_{3}$ and $\psi_{R}$ are as outlined previously in caption Fig. 3. Here the constraint $k_{1}+k_{2}+k_{3}=1$ has been relaxed to allow $k_{2}$ to vary. (d-e) Examples of inverted versions for the three main deformation modes with $\zeta^{\prime}(R)=1$. Blue (red) regions: negative (positive) Gaussian curvature. In all panels, $k_{B}=\sigma=1/10$.
In the bend dominated regime $k_{1}\!<\!k_{3}$ and for $\psi_{R}\!\approx\!\pi/2$, we observe pointy deformations on which the phase transitions from a vortex at the boundary to an aster at the center, Fig. 3e. In this configuration, the transition in the phase is localized to a ring of negative Gaussian curvature where bend distortions can be reduced at the cost of twist through the previously highlighted mechanism highlighted in Fig. 2d.
In the splay dominated regime $k_{1}\!>\!k_{3}$ and for $\psi_{R}\!\approx\!0$ we observe domed deformation on which the phase transitions from aster at the boundary to vortex at the center, Fig. 3d. Close to the boundary, the surface sharply deforms toward a cylindrical shape, abruptly reducing the splay at the cost of bend. On the cylindrical surface, the director field transitions from an aster to a vortex, further reducing splay and introducing additional twist and bend distortions. The surface flattens toward the top featuring a vortex. Overall, this buckling mode is able to achieve the highest degree of deformation.
Twist distortions are only possible in spiral regions of the defect when $\psi$ has an intermediate value, this is only observed in the deformation modes that feature a transition in the phase. Therefore the value of $k_{2}$ can be used to mediate the deformation of these surfaces. The splay dominated deformation mode, Fig. 3d, features twist on an extended cylindrical section of the surface. When $k_{2}$ is increased the cylindrical area is reduced and the spiral is pushed into the positive Gaussian curvature, Fig. 4a&c. This reduces the induced twist and the height of the deformation reduces asymptotically toward a finite value, cyan curve Fig. 4a. The radial transition in the phase of the bend dominated deformation, Fig. 3e, is entirely within the negative Gaussian curvature region, which introduces increased twist. When $k_{2}$ is increased, the only way to reduce the twist is to reduce the negative Gaussian curvature, which in turn reduces the magnitude of the deformation and eventually suppress this mode, magenta curve Fig. 4a. Conversely, the conical aster deformation, Fig. 3c, features no transition in the phase, hence no twist, and is approximately unaffected by increasing $k_{2}$, black curve Fig. 4a. In fact, when $k_{2}$ is reduced to near zero, the director will locally distort to reduce splay and introduce twist, Fig. 4b.
We additionally observe bistability for these solutions. All surfaces shown thus far are identical under the reflection $\zeta\rightarrow-\zeta$. This symmetry can be broken by setting the boundary conditions $\zeta^{\prime}(R)\neq 0$. All three previously identified buckling modes have a stable configuration in which $\zeta^{\prime}$ changes sign closer to the center suggesting a similar mechanism could play a role in driving invagination, see Figs. 4d-f.
Finally, this study reveals a novel mechanism for the spontaneous fusion of half-integer topological defects on fluid membranes. Consider the special case $\sigma=0$ and $k_{1}\!=\!k_{2}\!=\!k_{3}\!=\!k/3$. It is known that the energy minimizing nematic field with a total charge of $+1$ on a flat disc has two $+1/2$ defects. In this case, the energy scales, up to numerical pre-factors of order $1$, as ${\cal F}_{f}\sim\pi k/3\log(1/\Delta)$ [46]. However, the free-energy of a conical surface with an aster at the apex scales sub-linearly with $k$, (i.e ${\cal F}_{c}\sim\pi k_{B}\sqrt{(4k/3k_{B}-1)})\log(1/\Delta)$). Therefore, a critical threshold $(k/3k_{B})_{c}={\cal O}(1)$ arises from the balance between the energies of these two states. Thus, if $(k/3k_{B})>(k/3k_{B})_{c}$, the pair of $+1/2$ topological defects can spontaneously fuse by deforming the surface out of plane. Indeed, past works have shown that this process can occur for prescribed conical surfaces [25].
This mechanism for spontaneous deformation of fluid membranes can drive out-of-plane deformations of biological system, such as a film of cytoskeletal filaments on a supported lipid bilayer or a cell monolayer on a supported elastic substrate, see Sec. 6 in [44]. For instance, actin and microtubule films can exhibit nematic phases [47, 48, 49]. The measured reduced Frank constant is in the range of $k=10^{-8}-10^{-7}~{}\mu N\mu m$ for actin filament films [50, 51] and is $k=10^{-3}~{}\mu N\mu m$ for microtubule films [52]. Using typical values for the bending rigidity and surface tension of lipid bilayers and ignoring other possible elastic contributions, allows one to evaluate the proximity to the instability threshold (S5). In particular, in the bending dominate regime $k_{B}\gg\sigma R^{2}$, the critical threshold is of order of $(k/k_{B})_{c}\sim 1$, and the ratio is $k/k_{B}=1$ for a film of actin filaments and is $k/k_{B}=10^{4}-10^{5}$ for a film of microtubules. In the tension dominated regime $k_{B}\ll\sigma R^{2}$, the critical threshold is $(k/\sigma R^{2})_{c}\sim 1$ and therefore it can vary with the disc radius $R$. One can determine an upper bound by replacing the radius $R$ with the layer thickness $h$, which is the smallest lengthscale. This bound is $k/\sigma h^{2}<10$ for a film of both types of cytoskeletal filaments. These estimations suggest that in low-tension regimes, biological system can induce out-of-plane deformation by relieving elastic stresses at integer topological defects. In fact, this phenomenon may have been observed in Ref. [38] that studied a film of microtubules encapsulated on a vesicle. In this work, the authors reported that lowering the surface tension leads to spindle-like vesicles with two $+1$ defects localized at the spindle poles.
Shape dynamics driven by out-of-equilibrium processes could also be influenced by this mechanism. For instance, because the aster tends to facilitate surface deformations, it can be a nucleation point for out-of-plane deformations on active fluid membranes, like bacterial biofilms or cell monolayers [38, 37, 53, 40, 29].
Acknowledgements.
We are grateful to Isabelle Bonnet, Mathieu Dedenon, Karsten Kruse, Jean François Joanny, Jacques Prost, and Feng-Ching Tsai for insightful discussions. DJGP acknowledges funding from the Swiss National Science Foundation under starting grant TMSGI2 211367. This project has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 847718. Q.C acknowledges funding from Institut Curie EuReCa PhD Programme. This publication reflects only the author’s view and that the European Research Agency is not responsible for any use that may be made of the information it contains.
References cited in SI: [38, 54, 37, 39, 40, 55, 29, 47, 48, 49, 56, 57, 58, 59, 50, 51, 60, 61, 62, 52, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 46, 75, 76]